Syllabus UTK – M251 – Matrix Algebra Fall 2003, Jochen Denzler, MWF 1:25–2:15, Ayres 318

Textbook: Howard Anton: Elementary Linear Algebra — Odd enough the book has even answers to even numbered questions. But we’ll get around this oddity.

Course contents: Basically, Ch. 1–7 of the book, with a few ommissions, and a bit from Ch. 8, 9 blended in. See special notes for an overview.

Grade: We’ll have 3 in-class and one final exam, and homework. They count towards the grade as follows: either \frac{1}{7} each in-class exam, and \frac{2}{7} each for the final and for the homework, or else \frac{1}{6} each in-class exam, and the homework, and \frac{2}{6} for the final, whichever gives the better average for each individual.

Note: I will NOT give an extensive review before in-class exams. This would encourage ineffective learning strategies. Make sure to assemble the material in your brain as a coherent and meaningful entity, not as a bag of of single tricks and skills. That’s more rewarding and prepares your better, and you’ll need to relearn less before the comprehensive final.

Homework: Each homework problem will be graded with either 0, 1 or 2 points, unless otherwise specified. This puts the responsibility on you to hand in decent solutions and do calculations correctly. You will typically not be able to find solutions in the back of the book, because I will make up my own problems. However you will find very similar problems (with different numbers) in the book. You can use them as needed; I will not assign gazillions of cloned problems for training ad nauseam, such as to leave you time to think and breathe.

There will often be ways how you can check the correctness, or at least the plausibility of your result. And the question “how can I check my answer to this problem, once I have found an answer?”, is a good question to ponder about, or to ask me if you don’t see it.

I may occasionally assign true/false or conceptual questions from the book and then schedule a brief quiz with a selection, essentially, from these questions. A quiz question will then count as a homework question. But be sure that you understand, rather than memorize, the answers. I am determined and able to detect and defeat parroted answers. I am even more determined (and hopefully able;-) to help you in the office hour if you have difficulties in understanding the concepts (as indicated, e.g., by an urge to memorize).

I have scheduled regular office hours Monday 11:15-12:00 and Wednesday 2:30-3:30. You are welcome to request an appointment at other times or drop in. I accomodate drop-ins, whenever feasible, even though I cannot always guarantee immediate availability. My office is Ayres 317 E, phone 4-5325. Email is denzler@math.utk.edu, but I may not read it for half a day or for an entire weekend.